The men's world record for swimming 1500.0 m in a long course pool (as of 2007) is 14 min 34.56 s. At this rate, how many seconds would it take to swim 0.600

miles (1 mi = 1609 m)?


1
Expert's answer
2021-02-05T04:02:53-0500

At first, time should be expressed in seconds:

t1=(14×60)+34.56=874.56(s).

v=l/t=1500.0/874.56=1.77(m/s).

l2=0.600×1609=965.4(m)

t2=l2/v=965.4/1.77=545.42(s)=9 min 5.42 s